# A New Determination of the Mass of NGC 3603-A1: the Most Massive Binary Known?

### v1

@@ Abstract
The star NGC 3603-A1 has long been known to be a very massive binary, consisting of a pair of O2-3If*/WN5-6 stars, which show Wolf-Rayet-like emission due to their luminosities being near the Eddington limit. The system has been poorly characterized until now, due to the difficulties of obtaining reliable radial velocities from broad, blended emission lines and the extreme crowding in the cluster. However, previously unpublished archival HST/STIS spectra revealed that some of the upper Balmer lines (seen in absorption) are well separated at favorable orbital phases, prompting us to obtain our own carefully-timed new HST/STIS spectra, which we have analyzed along with the older data. Radial velocities measured from these spectra allow us to obtain an orbit for this 3.77298 day system. We also used archival STIS imaging of the cluster to obtain a more accurate light curve for this eclipsing system, which we then modeled, yielding the orbital inclination and providing values for the stellar radii and temperatures. Together, these data show that the NGC 3603-A1 system consists of a $93.3\pm 11.0M_{\odot}$ O3If*/WN6 primary with an effective temperature of 37,000 K, and a $70.4\pm 9.3M_{\odot}$ O3If*/WN5 secondary that is slightly hotter, 42,000 K. Although a more massive binary is known in the LMC, NGC 3603-A1 is the most massive binary in our own Galaxy for which a direct measurement of its mass has been made by a fundamental method. The secondary has been spun up by mass accretion from the primary, and we discuss the evolutionary status of this intriguing system.
@@ 4.1 Is NGC 3603-A1 “the most massive star known?”
At one time, NGC 3603-A1 was touted as ‘the most massive binary.” While this is not completely correct, it may not be completely wrong, either: we find that although there is a binary with more massive components known in the LMC, the mass of the primary in NGC 3603-A1 is the largest of any known star in the Milky Way that has a direct, Keplerian measurement.
@@
Thus, we conclude that NGC 3603-A1 is one of the most massive binaries known, with masses of $93.3\pm 11.0M_{\odot}$ and $70.4\pm 9.3M_{\odot}$ for the primary and secondary. The primary has the highest known mass of any star in the Milky Way that has a measurement based on Keplerian motion and a well-analyzed light curve. In Table 4.1 we list the spectral types and masses of the highest mass binaries with well-determined parameters.
@@
Mk 39 O2.5If*/WN6 + O3V-III $105\pm 11$ $80\pm 11$ 5
R144 WN5/6h + WN6/7h $74\pm 4$ $79\pm 4$ 6
R136-038 O3V + O6V $56.9\pm 0.6$ $23.4\pm 0.2$ 7
References. — 1–This paper; 2–A. Z. Bonanos et al. 2004; 3–G. Rauw et al. 2005; 4–M. E. Lohr et al. 2018; 5–A. M. T. Pollock et al. 2025; 6–T. Shenar et al. 2021; 7–P. Massey et al. 2002.
@@ 4.2 Evolutionary Status
One of the most interesting things, however, is that both the primary and secondary have optically thick winds with spectra dominated by broad WR-like emission. The luminosities at which stars develop such winds and form O2-3If*/WN5-6 stars has never been established observationally. This is true for the other high-mass binaries, Mk 39 and WR20a, as we see in Table 4.1. (We speculate as to whether the R144 components have been misclassified, and that these are also slash stars.) It is unclear whether this is purely due to their high luminosities, or if binary interactions are also contributing to the formation of optically thick winds.
@@
Lowell Observatory sits at the base of mountains sacred to tribes throughout the region. We honor their past, present, and future generations, who have lived here for millennia and will forever call this place home. This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute (STScI), which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. These observations are associated with programs GO-10602 (PI: Maiz Apellaniz), GO-11626 (PI: Massey), GO-12615 (PI: Schnurr), and GO-17527 (PI: Massey). We thank our HST/STIS program scientist Alex Fullerton and our program coordinator Christian Soto for their help throughout the process of obtaining our data. Support for programs GO-11626 and GO-17527 were provided to P.M. by NASA through a grant from the STScI. Support for S.B.’s participation via the National Science Foundation’s Research Experiences for Undergraduates was made possible through NSF awards 1852478 and 1950901. We thank Dr. Jan Eldridge for interesting correspondence on this system. We also thank Mike Fitzpatrick for help with several IRAF-related issues. NOIRLab IRAF is distributed by the Community Science and Data Center at NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the NSF. Assistance in Python programming (both for error propagation and disentanglement) was provided by ChatGPT-4 (OpenAI et al., 2023). An anonymous referee helped improve our presentation. This work also made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.

### v3

@@ Abstract
The star NGC 3603-A1 has long been known to be a very massive binary, consisting of a pair of O2-3If*/WN5-6 stars, which show Wolf-Rayet-like emission due to their luminosities being near the Eddington limit. The system has been poorly characterized until now, due to the difficulties of obtaining reliable radial velocities from broad, blended emission lines and the extreme crowding in the cluster. However, previously unpublished archival HST/STIS spectra revealed that some of the upper Balmer lines (seen in absorption) are well separated at favorable orbital phases, prompting us to obtain our own carefully-timed new HST/STIS spectra, which we have analyzed along with the older data. Radial velocities measured from these spectra allow us to obtain an orbit for this 3.77298 day system. We also used archival STIS imaging of the cluster to obtain a more accurate light curve for this eclipsing system, which we then modeled, yielding the orbital inclination and providing values for the stellar radii and temperatures. Together, these data show that the NGC 3603-A1 system consists of a $93.3\pm 11.0M_{\odot}$ O3If*/WN6 primary with an effective temperature of 37,000 K, and a $70.4\pm 9.3M_{\odot}$ O3If*/WN5 secondary that is slightly hotter, 42,000 K. Although a more massive binary is known in the LMC, NGC 3603-A1 is as massive as any binary known in our own Galaxy for which a direct measurement of its mass has been made. The secondary has been spun up by mass accretion from the primary, and we discuss the evolutionary status of this intriguing system.
@@ 4.1 Is NGC 3603-A1 “the most massive star known?”
At one time, NGC 3603-A1 was touted as ‘the most massive binary.” While this is not completely correct, it may not be completely wrong, either: we find that although there is a binary with more massive components known in the LMC, the mass of the primary in NGC 3603-A1 is as large as any known star in the Milky Way that has a direct, Keplerian measurement.
@@
Thus, we conclude that NGC 3603-A1 is one of the most massive binaries known, with masses of $93.3\pm 11.0M_{\odot}$ and $70.4\pm 9.3M_{\odot}$ for the primary and secondary. In the Milky Way, only WR21a is known to have an equally massive primary. In Table 4.1 we list the spectral types and masses of the highest mass binaries with well-determined parameters.
@@
Mk 39 O2.5If*/WN6 + O3V-III $105\pm 11$ $80\pm 11$ 6
R144 WN5/6h + WN6/7h $74\pm 4$ $79\pm 4$ 7
R136-038 O3V + O6V $56.9\pm 0.6$ $23.4\pm 0.2$ 8
References. — 1–This paper; 2–R. H. Barbá et al. 2022; 3–A. Z. Bonanos et al. 2004; 4–G. Rauw et al. 2005; 5–M. E. Lohr et al. 2018; 6–A. M. T. Pollock et al. 2025; 7–T. Shenar et al. 2021; 8–P. Massey et al. 2002.
@@ 4.2 Evolutionary Status
One of the most interesting things, however, is that both the primary and secondary have optically thick winds with spectra dominated by broad WR-like emission. The luminosities at which stars develop such winds and form O2-3If*/WN5-6 stars has never been established observationally. This is true for the other high-mass binaries, Mk 39, WR21a, and WR20a, as we see in Table 4.1. (We speculate as to whether the R144 components have been misclassified, and that these are also slash stars.) It is unclear whether this is purely due to their high luminosities, or if binary interactions are also contributing to the formation of optically thick winds.
@@
Lowell Observatory sits at the base of mountains sacred to tribes throughout the region. We honor their past, present, and future generations, who have lived here for millennia and will forever call this place home. This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute (STScI), which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. These observations are associated with programs GO-10602 (PI: Maiz Apellaniz), GO-11626 (PI: Massey), GO-12615 (PI: Schnurr), and GO-17527 (PI: Massey). We thank our HST/STIS program scientist Alex Fullerton and our program coordinator Christian Soto for their help throughout the process of obtaining our data. Support for programs GO-11626 and GO-17527 were provided to P.M. by NASA through a grant from the STScI. Support for S.B.’s participation via the National Science Foundation’s Research Experiences for Undergraduates was made possible through NSF awards 1852478 and 1950901. Additional support was also provided to P.M. through AST-2307594. We thank Dr. Jan Eldridge for interesting correspondence on this system. We also thank Mike Fitzpatrick for help with several IRAF-related issues. NOIRLab IRAF is distributed by the Community Science and Data Center at NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the NSF. Assistance in Python programming (both for error propagation and disentanglement) was provided by ChatGPT-4 (OpenAI et al., 2023). An anonymous referee helped improve our presentation. This work also made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.